OpenStudy (anonymous):

Solve -2(x - 1) = 16 Help Please

OpenStudy (jamesj):

First thing I would do: divide both sides by -2. What do you get with that one step?

OpenStudy (anonymous):

-3x -2 =16

OpenStudy (anonymous):

x=-8

OpenStudy (jamesj):

No, you're doing something else here and that's ok. You are expanding out the brackets. If you do that you get, after that one step: -2x -(-2) = 16 Hence -2x + 2 = 16 Can you solve it now?

OpenStudy (jamesj):

@RobZBHayes Your answer is wrong. And in any case, please do not ever give just an answer without showing working.

OpenStudy (anonymous):

ya thanks the answer should be -7

OpenStudy (jamesj):

Right.

OpenStudy (anonymous):

The first thing you would have to do is use the distributive property and multiply -2 by x and -2 by -1, that would give you -2x+2=16, then you would subtract 2 from sixteen giving you -2x=14, then you would divide 14 by -2 to give you x=-7. I hope that helps

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

Yes you're right @JamesJ, however if you look at my explanation I corrected my work and explained how to get there. Does that make you feel better?